OpenStudy (anonymous):

Write an equation in slope intercept form for the line containing the points (3, –17) and (–4, 18).

OpenStudy (anonymous):

slope = (18-(-17))/(-4-3) slope = 35/-7 slope = -5 y-intercept: x = 0 -17 + (3x5) -17 + 15 = -2 y = -5x-2

OpenStudy (anonymous):

Thank you i was stuck on that one.